OpenStudy (anonymous):

Someone please help: Solve x^2 - 3x = -8.

OpenStudy (anonymous):

first of all we will calculate the discriminant "d" of x^2-3x+8=0 d=(-3)^2-4*8*1=9-32=-23<0 so the equation has no solution in IR do you want the complex solutions ??

OpenStudy (anonymous):

How did you get those numbers?

OpenStudy (anonymous):

Oh nevermind.

OpenStudy (anonymous):

It says to solve using the quadratic formula.

OpenStudy (anonymous):

okay so I'm going to do it step by step

OpenStudy (anonymous):

I'm gonna start with the general case ax^2+bx+c=0 in our case a=1, b=-3 and c=8 the steps so that you can get the quadratic formula are: step 1: x^2 + \frac{b}{a} x= -\frac{c}{a}.

OpenStudy (anonymous):

I'm sorry the last expression is: x^2+b/a*x=-c/a so here we move c to the other side and we devide by a

OpenStudy (anonymous):

step 2 we add to our expression ((1/2)*(b/a))^2 to both sides of the equality x^2+b/a*x+((1/2)*(b/a))^2=-c/a+((1/2)*(b/a))^2

OpenStudy (anonymous):

on the first hand we have: x^2+b/a*x+((1/2)*(b/a))^2=(x+b/(2a))^2 (you can check it)

OpenStudy (anonymous):

and now all we have to do is to replace by the values of a, b and c in (x+b/(2a))^2=-c/a+((1/2)*(b/a))^2 a=1, b=-3 and c=8 (x-3/2)^2=(-16+9)/2)=-7/2 as you can observe (x-3/2)^2 is positive and it can't equal something negative conclusion: the equation has no solutions.

OpenStudy (anonymous):

The problem is asking for a specific answer using imaginary numbers: i

OpenStudy (anonymous):

So far I have: x = 3 +/- (sqrt)[-17] / 2

OpenStudy (anonymous):

okay you want complex solutions

OpenStudy (anonymous):

The answer includes imaginary number i

OpenStudy (anonymous):

so let's start from the last thing I wrote (x-3/2)^2=(-16+9)/2)=-7/2

OpenStudy (anonymous):

we can write (-7/2) = (i*sqrt(7/2))^2 right !

OpenStudy (anonymous):

Is that the simplified form?

OpenStudy (anonymous):

no not yet actually I made a mistake in calcula it's -7/4 so (-7/4)=(i*sqrt(7)/2)^2 we replace (x-3/2)^2=(i*sqrt(7)/2)^2

OpenStudy (anonymous):

Okay what's the next step?

OpenStudy (anonymous):

we pass to the sqrt of the expression and we get abs(x-3/2)=i*sqrt(7)/2

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

is it clear until now ?

OpenStudy (anonymous):

Mostly.

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

First move -8 to left hand side , for that you have to add 8 on both sides , then you will get quadratic equation . So use quadratic formula and solve for x ,.... Hope this will help you

OpenStudy (anonymous):

abs(x-3/2)=i*sqrt(7)/2 so we have two cases (|a|=b -->a=b or a=-b) so x-3/2=i*sqrt(7)/2 or x-3/2=-i*sqrt(7)/2 i.e. x=3/2+i*sqrt(7)/2 or x=3/2-i*sqrt(7)/2 those are the solutions

OpenStudy (anonymous):

Neither of those fit the answer choices.

OpenStudy (anonymous):

both of them are correct we don't have one solution but we have a set of solutions {3/2+i*sqrt(7)/2,3/2-i*sqrt(7)/2} but I recommand you to do again the calcul if there is any mistake

OpenStudy (anonymous):

all of the choices are either to the sqrt of 29 or 23